### v3

# Capacity Approximations for Insertion Channels with Small Insertion Probabilities

Abstract
Channels with synchronization errors, exhibiting deletion and insertion errors, find practical applications in DNA storage, data reconstruction, and various other domains. Presence of insertions and deletions render the channel with memory, complicating capacity analysis. For instance, despite the formulation of an independent and identically distributed (i.i.d.) deletion channel more than fifty years ago, and proof that the channel is information stable, hence its Shannon capacity exists, calculation of the capacity remained elusive. However, a relatively recent result establishes the capacity of the deletion channel in the asymptotic regime of small deletion probabilities by computing the dominant terms of the capacity expansion. This paper extends that result to binary insertion channels, determining the dominant terms of the channel capacity for small insertion probabilities and establishing capacity in this asymptotic regime. Specifically, we consider two i.i.d. insertion channel models: insertion channel with possible random bit insertions after every transmitted bit and the Gallager insertion model, for which a bit is replaced by two random bits with a certain probability. To prove our results, we build on methods used for the deletion channel, employing Bernoulli$(1/2)$ inputs for achievability and coupling this with a converse using stationary and ergodic processes as inputs, and show that the channel capacity differs only in the higher order terms from the achievable rates with i.i.d. inputs. The results, for instance, show that the capacity of the random insertion channel is higher than that of the Gallager insertion channel, and quantifies the difference in the asymptotic regime.
Index Terms:
  Channels with synchronization errors, insertion channel, Gallager insertion channel, channel capacity, information stability, deletion channel.
I Introduction
Synchronization-error channels, including deletion, repetition, and insertion channels, serve as effective models for a variety of realistic and practical scenarios. Such channel models find extensive applications in DNA coding [1, 2], DNA-based storage systems [3, 4], magnetic data storage systems [5], data reconstruction [6], and numerous other domains. Despite the wide range of application areas, the capacity of channels with synchronization errors remains an underexplored research area. In [7], Dobrushin establishes counterparts to Shannon’s discrete memoryless channel’s (DMC’s) capacity formulation [8] for a wide range of channels with synchronization errors. He demonstrates that the channel capacity for such channels can be defined by the maximal mutual information per bit and can be achieved through a random coding scheme. In [9, 10], the authors consider channels with synchronization errors governed by a Markov chain. It is shown that if the underlying Markov chain is stationary and ergodic, the information capacity exists and is equal to the coding capacity. This result generalizes Dobrushin’s coding theorem [7] on the information stability of channels with memoryless synchronization errors.
Despite the early result establishing the information stability of the memoryless synchronization error channels, the calculation of the exact capacity proved extremely challenging. For instance, even for the seemingly simple channel model of independent and identically distributed (i.i.d.) deletion channel, the channel capacity is unknown. As such the literature on channels with synchronization errors predominantly focuses on establishing lower and upper bounds for channel capacity. In [11], Gallager proposed using a pseudo-random string to convolutional codes with sequential decoding to correct synchronization and substitution errors. Tree codes are utilized in [12] to obtain lower bounds for insertion/deletion channels. In [13], a lower bound for the deletion channel is established by demonstrating a connection between deletion channels and the Poisson-repeat channel using jig-saw decoding. This lower bound is further improved in [14] for the capacity of channels with independent and identically distributed (i.i.d.) deletions and duplications.
To develop upper bounds on the capacity of channels with synchronization errors, typical technique is to employ a genie-aided approach with additional side information, which results in more tractable channel models. In [15], an upper bound on the deletion channel capacity is derived using undeletable markers. Similarly, Ref. [16] provides deletion channel upper bounds coupled with lower bounds, by using different genie-aided systems. In [17], an upper bound for the non-binary deletion channel is obtained. Furthermore, in [18], capacity upper bounds are obtained for more general synchronization error channels exhibiting insertions, deletions and substitutions. In [19], using a systematic approach based on convex programming and real analysis, the author derives upper bounds on the capacity of the binary deletion channel, channels with i.i.d. insertions and deletions, and the Poisson-repeat channel. This framework can be applied to obtain capacity upper bounds for any repetition distribution. A recent and comprehensive overview of the capacity results of synchronization-error channels is provided in [20].
Parallel to the results establishing upper and lower bounds on the channel capacity, there have also been efforts determining the behavior of the capacity of the deletion channel in the asymptotic regime of small deletion probabilities. For instance, [21] demonstrates that the capacity of the deletion channel is comparable to that of the binary symmetric error channel in the asymptotic case for small deletion probabilities. The authors establish that the trivial capacity lower bound for this channel, given by $1-H(p)$, is tight in the limit. They provide an upper bound of $1-(1-o(1))H(p)$, where the $o(1)$ term vanishes as the deletion probability $p\rightarrow 0$. More in line with our work in this paper, the works presented in [22, 23] leverage Dobrushin’s coding theorem [7] to approximate the capacity of the binary deletion channel using series expansions up to the second and third leading terms, respectively, for small deletion probabilities. These studies establish the channel’s capacity in the asymptotic regime of small deletion probabilities by calculating the dominant terms of the capacity expansion. Additionally, they demonstrate that the optimal input distribution can be obtained by perturbing the i.i.d. Bernoulli$(1/2)$ input distribution. In [24], a similar approach is applied to duplication channels, where the channel capacity is computed for small values of the duplication probability using a series expansion. Note that, although the duplication channel is closely related to the insertion channel and shares similarities with our channel model, it duplicates existing bits/symbols rather than inserting random bits/symbols, hence no runs are lost or no new runs are created, significantly simplifying the capacity calculation. Indeed, the capacity for the duplication channel can be computed with arbitrary prevision for any duplication probability [25].
In this paper, we adopt a similar systematic approach to insertion channels, which serve as practical models for certain real-world communication schemes including DNA storage. We consider two fundamental binary insertion channel models: 1) the simple insertion channel, in which a random bit is inserted after each bit with a certain probability, and 2) the Gallager insertion channel, in which each bit is replaced with two random bits with a certain probability. In both cases, we focus on small insertion probabilities. We explore the first two terms in the series expansion of these channel models’ capacities and establish their achievability using i.i.d. Bernoulli$(1/2)$ input distribution, and establish a matching converse for the dominant (lower order) terms.
Our main results are as follows: let $C(\alpha)$ be the capacity of an insertion channel with insertion probability $\alpha$. Then, for the simple insertion channel model with small $\alpha$ and any $\epsilon>0$, we have
[…]
where
$G_{1}=-$ $\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right).$
For the Gallager insertion channel, we obtain
[…]
with
$G_{2}=-$ $\log(e)-\frac{7}{8}+\frac{1}{4}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right),$
where $G_{1}\approx 0.4901$, and $G_{2}\approx-0.5865$, respectively. Approximations obtained by omitting the terms with higher order for these channel models are illustrated in Fig. 1. As may be anticipated, the Gallager insertion channel replaces a bit with two random bits; hence, for a specific input, the ambiguity of the channel output is higher compared to the simple insertion channel model with random bit insertion (without replacement). Therefore, intuitively, one would expect a lower rate for the Gallager insertion channel, which is verified by the results in Fig. 1.
[…]
Figure 1: Capacity approximation for both insertion channels.
The paper is organized as follows. Section II introduces the preliminaries and the channel models. The maximum rate for the described insertion channel models is decomposed into entropy terms in Section III. The capacity approximation for the simple insertion channel and the Gallager insertion channel are provided in Sections IV and V, respectively. The paper is concluded in Section VI.
Notation: Throughout the paper, we use base-two logarithms. Additionally, we note that as $\alpha$ approaches 0, we have
$\log(1-\alpha)\approx-\alpha\log(e),$
where $e$ represents Euler’s number.
II Preliminaries
We consider insertion channel $W_{n}$ which takes $n$-bit binary vector $X^{n}$ as an input and outputs $Y(X^{n})$. We note that the inserted bits are independent of both other insertions and the original input bits. Specifically, we focus on the following two specific channel models in this paper:
1) Simple Insertion Channel: There is a random bit insertion after each bit with probability $\alpha$. We define the insertion pattern for this channel model by vector $A^{n}\in\{0,1\}^{n}$ whose $i$-th element is defined by
$A_{i}=\begin{cases}1,&\text{if there is an insertion after the }i\text{-th bit of }X^{n}\\ 0,&\text{otherwise}\end{cases}$
for $i\in\{1,2,\cdots,n\}$. The corresponding inserted bit vector is denoted by $B^{n}\in\{0,1\}^{n}$ whose $i$-th element is
$B_{i}=\begin{cases}0,&\text{if }A_{i}=0\\ U_{i},&\text{if }A_{i}=1\end{cases}$
where $U_{i}$ are i.i.d. $\text{Ber}(1/2)$ random variables. Note that value of $B_{i}$ for $A_{i}=0$ is not important and it is arbitrarily set to 0.
2) Gallager Insertion Channel: Each bit is replaced with two independent random bits with probability $\alpha$ as described in [26]. The insertion pattern for this model is defined by $A^{n}\in\{0,1\}^{n}$ whose $i$-th element is
$A_{i}=\begin{cases}1,&\text{if the }i\text{-th bit of }X^{n}\text{ is replaced }\\ &\text{with two random bits}\\ 0,&\text{otherwise}\end{cases}$
for $i\in\{1,2,\cdots,n\}$. For this channel model, we define two insertion bit vectors as $B^{n}=(B^{1,n},B^{2,n})$ where $B^{j,n}\in\{0,1\}^{n}$ with $j\in\{1,2\}$ and their $i$-th element is defined similarly to that of the insertion channel model, i.e.,
$B^{j}_{i}=\begin{cases}0,&\text{if }A_{i}=0\\ U_{i,j},&\text{if }A_{i}=1\end{cases}$
for $j\in\{1,2\}$, where $U_{i,j}$ are i.i.d. and $\sim\text{Ber}(1/2)$.
Note that for both channel models, the expected length of the output vector $Y(X^{n})$ is $\mathbb{E}[n^{\prime}]=n+\alpha n$, and we will use the notation $Y^{n^{\prime}}$ for the output vector.
In [7], Dobrushin showed that the maximum rate for reliable communication can be formulated by the maximal mutual information per bit for memoryless channels with synchronization errors. It is also shown that information stability holds and Shannon capacity exists. This approach can be used to express the capacity of the insertion channels above, as stated in the following theorem.
Theorem 1.
(Taken from [7]) For the memoryless channels with finite expectation of output vector length, consider the definition
$C_{n}\equiv\frac{1}{n}\max_{p_{X^{n}}}I(X^{n};Y(X^{n})).$ (4)
Then, the following limit exists and defines the capacity of the channels with synchronization errors
$C\equiv\lim_{n\rightarrow\infty}C_{n}=\inf_{n\geq 1}C_{n}.$ (5)
Furthermore, as stated in [7, Th. 5], the input $(X_{1},X_{2},\cdots,X_{n})$ can be considered as consecutive coordinates of a stationary ergodic process as in the following lemma.
Lemma 2.
(Taken from [23]) With $X_{i}\in\{0,1\}$ and stationary and ergodic process $\mathbb{X}=\{X_{i}\}_{i\in\mathbb{Z}}$, the limit $I(\mathbb{X})=\lim_{n\rightarrow\infty}\frac{1}{n}I(X^{n};Y(X^{n}))$ exists and we have
$C=\sup_{\mathbb{X}\in{\mathcal{S}}}I(\mathbb{X}),$ (6)
where $\mathcal{S}$ denotes the class of stationary and ergodic processes that take binary values.
See [23] for the proofs of Theorem 1 and Lemma 2.
We define the rate achieved by the stationary and ergodic process $\mathbb{X}$ as:
$I(\mathbb{X})=\lim_{n\rightarrow\infty}\frac{1}{n}I(X^{n};Y(X^{n})).$ (7)
Throughout this paper, we will use the following definition: for the binary process $\mathbb{X}$, the consecutive blocks of 0’s and 1’s in $X=(X_{1},X_{2},\cdots,X_{n})$ are called runs, and we investigate the capacity of insertion channels with the stationary ergodic input process $\mathbb{X}$, denoted by $\mathcal{S}$.
We are interested in studying the capacity of insertion channels for small insertion probabilities. Hence, we focus on an insertion channel with insertion probability $\alpha=0$ (namely, a perfect channel), and note that the capacity of such a channel is 1, and it is achieved with i.i.d. Bernoulli($1/2$) input distribution. We expect that if we consider i.i.d. input distributions for insertion channels with small insertion probabilities, the achievable rate will not deviate from the channel capacity by much. In fact, we will obtain our achievability results by using i.i.d. Bernoulli($1/2$) inputs, and by proving that the rate achieved matches the first two order terms of the actual channel capacity (by coupling the results with relevant converse bounds).
To prove the converse results, taking a run-length perspective, we will assume that no run exceeds a certain length. We first will show that limiting the run length does not lead to significant loss if the threshold is sufficiently large. Then, we will derive an upper bound on $I(\mathbb{X})$ over all stationary and ergodic processes. The obtained upper bound will match the achievable rates (with i.i.d. inputs) for the two orders terms of the channel capacity, hence proving the channel capacity approximation for the insertion channel model for small insertion probabilities.
III Decomposing Rate: Understanding $I(X^{n};Y(X^{n}))$
In this section, we identify the different terms necessary to determine the maximum rate for both insertion channel models under consideration. First, we define the runs for the input bit vector, which consists of consecutive bits with the same polarity, and a corresponding vector representing the lengths of the output bits (including the inserted bits). Using the channel models, insertion pattern, and inserted bit vector defined in Section II, we analyze the contributing factors to the capacity of the simple and Gallager insertion channels, which will be computed in detail in subsequent sections.
We use $X(j)$ to represent the $j$-th run in the input vector $X^{n}$, while $Y(j)$ denotes the corresponding sequence of bits in the output vector $Y^{n^{\prime}}$. We define $K$ as the vector of $|Y(j)|$’s, where ||$\cdot$|| ⋅ | indicates the number of bits in $Y(j)$. Assuming that there are $M$ runs in $X(j)$, one can express this as:
$X^{n}$ $=X(1)\dots X(M-1)X(M),$ (8)
$Y^{n^{\prime}}$ $=Y(1)\dots Y(M-1)Y(M),$ (9)
$K$ $=(|Y(1)|,\dots,|Y(M-1)|,|Y(M)|).$ (10)
We have the following expression for the rate given in (7) in terms of entropy:
$I(X^{n};Y)$ $=H(Y)-H(Y|X^{n})$ (11)
=$\stackrel{{\scriptstyle r_{1}}}{{=}}H(Y)-H(Y,K|X^{n})+H(K|X^{n},Y),$ (12)
where $(r_{1})$ follows from the chain rule for entropy. For the term $H(Y,K|X^{n})$, we have
$H(Y,K|X^{n})$ =$\stackrel{{\scriptstyle r_{2}}}{{=}}H(Y,K,A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{3}}}{{=}}H(Y,K|X^{n},A^{n},B^{n})+H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{4}}}{{=}}H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{5}}}{{=}}H(A^{n},B^{n})-H(A^{n},B^{n}|X^{n},Y,K),$ (13)
where $(r_{2})$ and $(r_{3})$ follow from the chain rule for entropy, $(r_{4})$ since $H(Y,K|X^{n},A^{n},B^{n})=0$, and $(r_{5})$ due to the independence of $A^{n}$ and $B^{n}$ with the input $X^{n}$. Consequently, we obtain the expression:
$I(X^{n};Y)=$ $H(Y)-H(A^{n},B^{n})+H(A^{n},B^{n}|X^{n},Y,K)+H(K|X^{n},Y).$ (14)
In this expression, $H(Y)$ quantifies the randomness in the output for all possible input sequences. $H(A^{n},B^{n})$ represents the uncertainty due to the insertion pattern and inserted bits, while $H(A^{n},B^{n}|X^{n},Y,K)$ indicates the ambiguity of the insertion pattern and the inserted bits after observing the input and output sequences, and run lengths. The last term, $H(K|X^{n},Y)$, captures the unpredictability of the run lengths after observing the input and output.
IV Capacity Approximation for the Simple Insertion Channel Model
In this section, we provide a capacity approximation for the first insertion channel model (dubbed the simple insertion channel) by computing the first two terms of the series expansion for small insertion probabilities. Our main result is given as follows.
Theorem 3.
Let $C(\alpha)$ be the capacity of the simple insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon})$ (15)
where
$G_{1}$ $=-\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right),$ (16)
which is approximated as $G_{1}\approx 0.4901$.
To establish this approximation, the first step involves calculating the terms in $I(X^{n},Y(X^{n}))$, as explicitly given in Section III and provided in (14). We elaborate on these calculations in the subsequent subsections.
IV-A Calculation of $H(A^{n},B^{n})$
We define the process denoting the insertion patterns corresponding to the sequence of insertion channels as $\mathbb{A}$, which is an i.i.d. Bernoulli($\alpha$) process that determines the locations of insertions, independent of the input. With this process and channel model, we represent the corresponding $n$-bit binary vector as $A^{n}$. Each element of $A^{n}$, denoted as $A_{i}$, is set to ‘1’ if there is an insertion after the $i$-th bit of the input vector, and ‘0’ otherwise. For the inserted bits, as described in Section II, $B^{i}=0$ if there is no insertion after the $i$-th bit while $B_{i}\sim\text{Ber}(1/2)$ otherwise.
Lemma 4.
We have
$\frac{1}{n}H(A^{n},B^{n})=h(\alpha)+\alpha,$ (17)
where $h(\cdot)$ is the binary entropy function.
Proof.
To compute $H(A^{n},B^{n})$ in (14), one can use the independence of the bits:
$H(A^{n},B^{n})=\sum_{i=1}^{n}H(A_{i},B_{i}).$ (18)
Note that the joint probability mass function (PMF) of $A_{i}$ and $B_{i}$ can be expressed as:
$P(A_{i}=a,B_{i}=b)=\begin{cases}1-\alpha,&\text{if }(a,b)=(0,0)\\ \frac{\alpha}{2},&\text{if }(a,b)=(1,0)\\ &\text{or }(a,b)=(1,1)\\ 0,&\text{otherwise}.\end{cases}$ (19)
As a result, we have
$H(A_{i},B_{i})$ $=-(1-\alpha)\log(1-\alpha)-\alpha\log(\alpha/2)$
$=-(1-\alpha)\log(1-\alpha)-\alpha\log(\alpha)+\alpha$
$=h(\alpha)+\alpha,$ (20)
where $h(\alpha)$ is the binary entropy function. We can also write
$H(A^{n},B^{n})=nh(\alpha)+n\alpha,$ (21)
hence the result follows. ∎
IV-B Calculation of $H(A^{n},B^{n}|X^{n},Y,K)$
To calculate $H(A^{n},B^{n}|X^{n},Y,K)$, we use an extended run perspective, where each bit at the two ends of the run is included in the run. Specifically, for a run of 0’s, the last bit of the left neighboring run and the first bit of the right neighboring run are considered part of the extended run; the same applies to runs of 1’s. First, we introduce a modified insertion process, defined by the modified vectors $\hat{A}^{n}$ and $\hat{B}^{n}$. In this modified process, we consider at most a single bit insertion for each extended run. If there is more than one insertion in an extended run, all the insertions in that run (not the extended run) are reversed (i.e., the run is transmitted without any insertion, and the input and output runs are identical). Note that for small insertion probabilities, typical insertion patterns will have insertions that are sufficiently spaced apart and will contribute to the higher-order terms in the capacity. Multiple insertions within a run will not affect the higher-order terms as will be shown later. Hence, one can obtain an accurate estimation of $H(A^{n},B^{n}|X^{n},Y,K)$ when the insertion probability is small (and consequently the capacity).
We further define stationary processes ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ to keep track of reversed insertions as follows:
${\mathbb{Z}}^{n}=\hat{\mathbb{A}}^{n}\oplus\mathbb{A}^{n},$ (22)
and
${\mathbb{V}}^{n}=\hat{\mathbb{B}}^{n}\oplus\mathbb{B}^{n}.$ (23)
With the modified insertion process, since we leave the runs unchanged if there are no insertions or exactly one insertion in the corresponding extended run, for these runs, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ consist entirely of 0’s. For the remaining runs, the corresponding segments in ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ are the same as in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$, respectively. Therefore, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ represent the segments in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$ that are reversed due to the modified insertion process, while the segments corresponding to extended runs with at most one insertion remain unchanged.
In the rest of this section, we first derive a bound on the difference between $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$ and $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$ for the modified insertion process and the original insertion process where $\hat{Y}$ is the output vector when the modified insertion process is applied, and $\hat{K}$ represents the run lengths of the output (defined similarly to (10), including both the bits in the original input and the inserted bits from the modified insertion process). This represents the approximation error incurred by assuming the modified insertion process instead of the original one. Later, we will show that this bound does not affect the first two-order terms, thereby not impacting our capacity approximation. Next, we compute $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$, which represents the ambiguity in the insertion pattern and inserted bits given the input, output, and associated run vector when the modified insertion process is applied. By combining these two steps, we will arrive at an approximation for $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$.
We begin by examining the difference between $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$ and $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$ for the modified and original insertion processes considering the following pairs:
$\left((X^{n},Y,K,A^{n},B^{n}),(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n})\right),$ (24)
and
$\left((X^{n},Y,K),(X^{n},\hat{Y},\hat{K})\right).$ (25)
Both of these pairs are in the form of $(T,E)$ such that $T$ is a function of $(E,Z^{n},V^{n})$ and $E$ is a function of $(T,Z^{n},V^{n})$. From the chain rule of entropy, we obtain:
$H(T)$ $=H(T,E)-H(E|T),$ (26)
$H(E)$ $=H(T,E)-H(T|E).$ (27)
Taking the difference, we get
$H(T)-H(E)=H(T|E)-H(E|T).$ (28)
By the chain rule, we can also write
$H(T|E)=H(T|E,Z^{n},V^{n})+$ $H(Z^{n},V^{n}|E)-H(Z^{n},V^{n}|E,T),$ (29)
Since $T$ can be written as a function of $(E,Z^{n},V^{n})$, we have $H(T|E,Z^{n},V^{n})=0$, and
$H(T|E)\leq H(Z^{n},V^{n}).$ (30)
Similarly, one can follow the same steps for $H(E|T)$ resulting in the same upper bound. Thus, we have:
$|H(T)-H(E)|\leq H(Z^{n},V^{n}).$ (31)
Therefore, we have:
$|H(X^{n},Y,K,A^{n},B^{n})-H(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n})|\leq H(Z^{n},V^{n})$ (32)
For the second pair, we also have:
$|H(X^{n},Y,K)-H(X^{n},\hat{Y},\hat{K})|\leq H(Z^{n},V^{n}).$ (33)
Combining (32) and (33) and using the definition of conditional entropy, we have:
$|H(A^{n},B^{n}|X^{n},Y,K)-H(\hat{A}^{n},\hat{B}^{n}$ $|X^{n},\hat{Y},\hat{K})|\leq 2H(Z^{n},V^{n}).$ (34)
Clearly, the processes $\mathbb{Z}$ and $\mathbb{V}$ are stationary, and we have $2H(Z^{n},V^{n})/n\leq 2h(z,v)$. Let us calculate the PMF $P(z=a,v=b)$.
$P(z=0,v=0)$ $=P(z=0)P(v=0|z=0)=P(z=0),$
$P(z=0,v=1)$ $=P(z=0)P(v=1|z=0)=0,$
$P(z=1,v=0)$ $=P(z=1)P(v=0|z=1)=\frac{1}{2}P(z=1),$
$P(z=1,v=1)$ $=P(z=1)P(v=1|z=1)=\frac{1}{2}P(z=1).$
Without loss of generality, let us focus on the first bit of a run. If there is an insertion at this bit, it will be reversed if at least one more insertion occurs within the extended run containing this run. For simplicity, we can consider the first run (the run containing the bit at position 0), which has a length of $L_{0}$ and extended run length $L_{0}+2$. Conditioned on $L_{0}$, the probability that $\{z=1\}$ is given by $\alpha\left(1-(1-\alpha)^{L_{0}+1}\right)$. Therefore, we can write
$P(z=1)=\alpha\left(1-\mathbb{E}[(1-\alpha)^{L_{0}+1}]\right).$ (35)
Thus, the joint PMF $P(z=a,v=b)$ can be written as:
$P(z=a,v=b)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}],\\ &\qquad\qquad\text{if }(a,b)=(0,0)\\ &\frac{1}{2}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}]\right),\\ &\qquad\text{if }(a,b)=(1,0)\text{ or }(1,1)\\ &0,\qquad\qquad\text{otherwise},\end{cases}$ (36)
We can use this expression to compute $h(z,v)$.
We further have $(1-\alpha)^{L_{0}+1}=1-\alpha(L_{0}+1)+\mathcal{O}(\alpha^{2})$. Hence, the PMF can be rewritten as:
$P$ $(z=a,v=b)$
=$=\begin{cases}1-(\mathbb{E}[L_{0}]+1)\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(0,0)\\ \frac{1}{2}\mathbb{E}[L_{0}+1]\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(1,0)\\ &\text{or }(a,b)=(1,1)\\ 0,&\text{otherwise}.\end{cases}$ (37)
One can upper bound the nonzero probabilities in (IV-B) as
$P(z=0,v=0)$ $\leq 1-2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (38)
$P(z=1,v=0)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (39)
$P(z=1,v=1)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (40)
which will be used to calculate the dominant terms of $h(z,v)$. This term identifies the loss incurred when restricting ourselves to the modified insertion process, i.e., when we assume there is at most one insertion in an extended run. Intuitively, these probabilities will not contribute to the dominant terms of the capacity expression. This will be made precise by working out the converse.
As mentioned earlier, to prove the achievability result, we will assume an i.i.d. Bernoulli(1/2) input distribution. Hence, the expectation of $L_{0}$ will be computed using
$p_{L_{0}}(l)=2^{-l},$ (41)
which gives $\mathbb{E}[L_{0}]=2$.
For the converse, additional work is required, which will be addressed in the following sections. Initially, we restrict the input run lengths to a specified threshold and show that this constraint does not impact the lower-order terms of the capacity. By deriving an upper bound for the converse and demonstrating its agreement with the achievability results for the first two-order terms, we will finalize the proof for the capacity approximation in the regime of small insertion probabilities.
Lemma 5.
Consider any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}H({A}^{n},{B}^{n}|X^{n},{Y}(X^{n}),K)=\frac{1}{2}\alpha\mathbb{E}[\log(L_{0})]-\eta,$ (42)
where $-2h(z,v)\leq\eta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v)$.
Proof.
The proof follows from (34) and Lemma 6 given next. ∎
Lemma 6.
Consider any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y}(X^{n}),\hat{K})=\frac{1}{2}\alpha\mathbb{E}[\log(L_{0})]-\delta,$ (43)
where $0\leq\delta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$.
Proof.
Having information $\hat{K}$ is equivalent to having a marker vector for the last positions of each run. Note that these indices are determined using the positions of the bits in $\hat{Y}(X^{n})$ corresponding to the last bit of a run (including the position with an insertion after the last bit if there is such an insertion) in $X$.
The proof consists of the following steps. First, we fix the channel input $x^{n}$ and any possible output $\hat{y}(x^{n})$. Then, we estimate (the $\log$ of) the number of possible realizations of $\hat{A}^{n}$ and $\hat{B}^{n}$ that might lead to the input/output pair $(x^{n},\hat{y})$ given the marker position vector $\hat{k}$. The final step is to take the expectation over $(x^{n},\hat{y},\hat{k})$.
Proceeding from left to right, and using the constraints on both $\hat{\mathbb{A}}^{n}$ and $\hat{\mathbb{B}}^{n}$, one can map each run of $\hat{y}$ to a run of $x^{n}$ given the marker vector $\hat{k}$. To simplify the exposition, let us focus on the $r$-th run of $x^{n}$ which consists of ‘0’s. $l_{r}$ denotes the length of this run (without the extended bits).
[…]
For this example, there are three cases:
  1. 1.
    No insertion: Since there is no insertion, there is a one-to-one mapping between the run of $x^{n}$ and $\hat{y}$. Hence, there is no ambiguity and no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
  2. 2.
    Insert bit ‘0’: There are $l_{r}$ possible locations (each of them being equally likely) for insertion in $\hat{A}$ given the marker vector $\hat{K}$.
    There is no additional ambiguity due to $\hat{B}$ (the corresponding bits of $\hat{B}$ will be all zero regardless of $l_{r}$ possible insertion positions).
    Hence, there will be a contribution of $\log(l_{r})$ to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
  3. 3.
    Insert bit ‘1’: Since the original run of $x^{n}$ consists of ‘0’s, inserting ‘1’ will reveal all the information regarding $\hat{A}$ and $\hat{B}$ given the marker vector $k$. Hence, there is no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
Note that, a similar argument can be made for other runs of ‘0’s or ‘1’s.
The total contributions due to case 2 previously explained (insertion of ‘0’ into runs of 0’s, and insertion of ‘1’ into runs of 1’s) can be expressed as a summation over all bits in the input for the stationary and ergodic process and given by:
$\frac{1}{2}\sum_{i=1}^{n}\hat{A}_{i}\log(l_{(i)}),$ (45)
where $l_{(i)}$ is the length of the run containing the $i$-th bit of $x^{n}$ and $\hat{A}_{i}$ is the $i$-th bit of the modified insertion pattern. Note that if there is no insertion for a certain bit, $\hat{A}_{i}=0$ and there will be no contribution to the summation.
We further have
[…]
$=(1-\alpha)^{l_{r}+1}\alpha.$ (46)
Hence,
$P(\hat{A}_{i}=1)=(1-\alpha)^{l_{(i)}+1}\alpha,$ (47)
and we can write $P(\hat{A}_{i}=1)\in(\alpha-(l_{(i)}+1)\alpha^{2},\alpha)$. Since, $l_{(i)}\geq 1$ and $l_{(i)}+1\leq 2l_{(i)}$, we can also write
$P(\hat{A}_{i}=1)\in(\alpha-2l_{(i)}\alpha^{2},\alpha).$ (48)
Inserting (48) into (45), taking expectation and letting $n\rightarrow\infty$, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})=\frac{1}{2}\alpha\mathbb{E}[\log(L_{0})]-\delta,$ (49)
where $0\leq\delta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$, which completes the proof of Lemma 6. ∎
IV-C Calculation of $H(K|X^{n},Y)$
Lemma 7.
For a simple insertion channel model with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(K|X^{n},Y)=\frac{\alpha}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}$ $(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right)+\epsilon_{1},$ (50)
where
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (51)
with $\epsilon>0$.
Proof.
To compute $H(K|X^{n},Y)$, we begin by defining a perturbed insertion process, denoted as $\check{\mathbb{A}}$. For the input process $\mathbb{X}$, we introduce $\check{\mathbb{Z}}$, whose elements correspond to two consecutive input runs with at most one insertion, equated to ‘0’. Otherwise, the elements mirror those of the original insertion process ${\mathbb{A}}$. To clarify further, for two consecutive input runs $\{S_{i},S_{i+1}\}$, we define ˇ$\check{\mathbb{Z}}^{i}$ as a binary process with all zeros when $\{S_{i},S_{i+1}\}$ has at most one insertion. Otherwise, $\check{Z}_{l}^{i}=1$ if $X_{l}\in S_{i}$ and $A_{l}=1$, $\forall i$. Utilizing the definition of $\check{Z}_{l}^{i}$, we further define $\check{\mathbb{Z}}$ in the following way:
$\check{Z}_{l}=\begin{cases}1,&\text{if }\exists i\text{ such that}\check{Z}_{l}^{i}=1\\ 0,&\text{otherwise.}\end{cases}$ (52)
This perturbed process yields a new insertion pattern defined as:
$\check{\mathbb{A}}\equiv\mathbb{A}\oplus\check{\mathbb{Z}}.$ (53)
A sample of the runs between $i$ and $i+4$ is provided in Table I, showing the changes in the perturbed insertion process compared to the original insertion process.
TABLE I: An example illustrating the input runs alongside the original and perturbed insertion processes.
run index: $i$ $i+1$ $i+2$ $i+3$ $i+4$
$S$ 000 $1$ 00 111 $0$
$A$ 010 1000 00 001 $0$
$\check{Z}$ 010 $0$ 00 000 $0$
$\check{A}$ 000 1000 00 001 $0$
Regarding the term $H(K|X^{n},Y)$, we first calculate the entropy using the perturbed process, followed by bounding the difference between the original insertion process and the perturbed process.
The output of this perturbed process is denoted by $\check{Y}(X^{n})$, and we define $\check{K}$ as follows:
$\check{K}=(|\check{Y}(1)|,\dots,|\check{Y}(M-1)|,|\check{Y}(M)|),$ (54)
where $\check{Y}(j)$ denotes the bits in the output corresponding to the $j$-th run of the input (including the original bits and insertions due to the perturbed process). Applying the chain rule, we express $H(\check{K}|X^{n},\check{Y})$ as:
$H(\check{K}|X^{n},\check{Y})$ $=H\left(|\check{Y}(1)|,\dots,|\check{Y}(M-1)|,|\check{Y}(M)|\Big{|}X^{n},\check{Y}\right)$
$=\sum_{j=1}^{M}H\bigg{(}|\check{Y}(j)|\Big{|}|\check{Y}(1)|,\dots,|\check{Y}(j-1)|,$
$\qquad X(1)X(2)\cdots X(M),\check{Y}(1)\check{Y}(2)\cdots\check{Y}(M)\bigg{)},$ (55)
where $X(1)X(2)\cdots X(M)$ and $\check{Y}(1)\check{Y}(2)\cdots\check{Y}(M)$ represent concatenations of $X(j)$’s and $\check{Y}(j)$’s, respectively. Therefore, $X^{n}=X(1)X(2)\cdots X(M)$ and $\check{Y}=\check{Y}(1)\check{Y}(2)\cdots\check{Y}(M)$.
In (IV-C), the condition on $|\check{Y}(1)|,\dots,|\check{Y}(j-1)|$, $X(1)X(2)\cdots X(j-1)$, and $\check{Y}(1)\check{Y}(2)\cdots\check{Y}(j-1)$ can be omitted. Hence, equivalently, we obtain:
$|\displaystyle H(\check{K}|X^{n},\check{Y})=\sum_{j=1}^{M}H\bigg{(}|\check{Y}(j)|$$\Big{|}X(j)\cdots X(M),\check{Y}(j)\cdots\check{Y}(M)\bigg{)}.$ (56)
The proof is completed by computing $H(\check{K}|X^{n},\check{Y})$ in (56) and establishing a bound for $|H({K}|X^{n},{Y})-H(\check{K}|X^{n},\check{Y})|$, as given in the Lemmas 8 and 9, respectively.
Lemma 8.
For a simple insertion channel with the perturbed insertion process and a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(\check{K}|X^{n},\check{Y})=\frac{\alpha}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)-\epsilon_{2}+\epsilon_{3},$ (57)
which is approximately $\approx 1.2885\alpha-\epsilon_{2}+\epsilon_{3}$ where
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right),$ (58)
and
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (59)
Proof.
We define:
$t_{j}\equiv H\left(|\check{Y}(j)|\Big{|}X(j)\cdots X(M),\check{Y}(j)\cdots\check{Y}(M)\right)$ (60)
With a slight change in notation, let us define $\bar{Y}(j^{\prime})$ as the runs in $\check{Y}$, meaning that $\bar{Y}(j^{\prime})$ is directly determined by the bits of $\check{Y}$. Consequently, we can rewrite $t_{j}$ as:
$t_{j}\equiv H\left(|\check{Y}(j)|\Big{|}X(j)\cdots X(M),\bar{Y}(j^{\prime})\cdots\bar{Y}(M^{\prime})\right).$ (61)
We define the length of $\bar{Y}(j)$ as $L_{j}$, denoted by $|\bar{Y}(j)|=L_{j}$. Consider four cases:
1. When $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: In this scenario, there is no insertion in $X(j)$. Thus, $t_{j}=0$, since we have sufficient information to determine $|\check{Y}(j)|$, given $X(j)\cdots X(M)$.
2. When $|\bar{Y}(j)|<|X(j)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the opposite polarity, for example, inserting a ‘0’ bit into a run of ‘1’s.
We can directly determine $|\check{Y}(j)|$ using $\bar{Y}(j)$, $\bar{Y}(j+1)$, and $\bar{Y}(j+2)$. Hence, $t_{j}=0$.
Note that $|\bar{Y}(j+1)|=1$, and the corresponding bit has the opposite polarity to $\bar{Y}(j)$ and $\bar{Y}(j+2)$.
3. When $|\bar{Y}(j)|>|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the same polarity, for example, inserting a ‘0’ bit into a run of ‘0’s.
Similar to the previous case, we can directly determine $|\check{Y}(j)|$ and $t_{j}=0$.
4. When $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|+1$: In this case, a single insertion occurs in $X(j)X(j+1)$, and we cannot distinguish between two different scenarios. Consider the following example:
$X(j)X(j+1)$ =[1 1 1 0 0 0 0]absentdelimited-[]$1$\ 1\ 0\ 0\ 0\ 0]= [ 1 1 1 0 0 0 0 ] (62)
$\bar{Y}(j)\bar{Y}(j+1)$ =[1 1 1 0 0 0 0 0]absentdelimited-[]$1$\ 1\ 0\ 0\ 0\ 0\ 0]= [ 1 1 1 0 0 0 0 0 ] (63)
Here, we cannot determine whether the ‘0’ bit is inserted at the end of $X(j)$ (the last possible position for insertion in $X(j)$) or at any of the possible $|X(j+1)|$ locations in $X(j+1)$. This scenario can be covered by the following two cases:
  • Case $\mathcal{V}_{1}$: If the ‘0’ bit is inserted at the end of $X(j)$:
    $|\check{Y}(j)|$ $=|\bar{Y}(j)|+1=|X(j)|+1$ (64)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j+1)|-1=|X(j+1)|$ (65)
    This case has a probability:
    $P(\mathcal{V}_{1})=\frac{1}{2}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}.$ (66)
  • Case $\mathcal{V}_{2}$: If the ‘0’ bit is inserted at any of the possible $|X(j+1)|$ locations in $X(j+1)$,
    $|\check{Y}(j)|$ $=|\bar{Y}(j)|=|X(j)|$ (67)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j+1)|+1=|X(j+1)|+1$ (68)
    This case has a probability:
    $P(\mathcal{V}_{2})=\frac{1}{2}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}|X(j+1)|.$ (69)
Combining both cases for $|\check{Y}(j)|=|X(j)|$ and $|\check{Y}(j)|=|X(j)|+1$, and defining $r_{j}\equiv|X(j)|$, we can write:
$|\check{Y}(j)|=\begin{cases}r_{j}+1,&\text{with probability }\frac{1}{2}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}\\ r_{j},&\text{with probability }\frac{1}{2}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}r_{j+1}\end{cases}$ (70)
Note that the only ambiguity in $|\check{Y}(j)|$ arises from (70), as all the previously described scenarios provide the necessary information for correct estimation. Hence, conditioned on case 4, we can normalize the probabilities as follows:
$|\check{Y}(j)|=\begin{cases}r_{j}+1,&\text{with probability }\frac{1}{r_{j+1}+1}\\ r_{j},&\text{with probability }\frac{r_{j+1}}{r_{j+1}+1}\end{cases}$ (71)
Consequently, we have:
$t_{j}=h\left(\frac{1}{r_{j+1}+1}\right).$ (72)
Thus, the expected contribution of this term to the sum is:
$\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $\frac{1}{2}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}(r_{j+1}+1)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{1}{r_{j+1}+1}\right),$ (73)
where $p_{L(k)}(l_{1},\dots,l_{k})$ is the joint distribution of input run lengths for $k$ runs.
It is worth noting that [^1] We have $(1-\alpha)^{x}\geq 1-\alpha x$ for $\alpha\rightarrow 0$ and $x\geq 1$. $(1-\alpha)^{r_{j}+r_{j+1}-1}\in\left(1-\alpha(r_{j}+r_{j+1}),1\right)$, hence (73) can be rewritten as:
$\frac{\alpha}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j+1}+1)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{1}{r_{j+1}+1}\right)-\epsilon_{2},$ (74)
with
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right),$ (75)
which is obtained by using the upper bound $p_{L(2)}(r_{j},r_{j+1})\leq 1$ in (75).
Note that in (74), the optimal distribution of the input run lengths $p_{L(2)}$ is used. To establish a bound for this term, as in [23, Lemma V.5], let us assume $p^{*}_{L(k)}(l_{1},\dots,l_{k})=2^{-\sum_{i=1}^{k}l_{i}}$. Consider any positive integer $k$, $\kappa<\infty$, and $H(\mathbb{X})>1-\alpha^{\gamma}$ with $\gamma>1/2$. Then, [23, Lemma V.5] implies
$\mid p_{L(k)}(l_{1},\dots,l_{k})-p^{*}_{L(k)}(l_{1},\dots,l_{k})\mid\leq\kappa\sqrt{k}\alpha^{\gamma/2}.$ (76)
Inserting (76) to the first term of the (74), one can write
$H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (77)
where
$A_{1}=\frac{1}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right),$ (78)
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (79)
and $\epsilon_{2}$’s range is defined in (75).
Note that we have $A_{1}\approx 1.2885$. ∎
Lemma 9.
For a simple insertion channel model with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$; the difference between $H(K|X^{n},Y)$ and $H(\check{K}|X^{n},\check{Y})$, corresponding to the original insertion and perturbed insertion processes, respectively, satisfies:
$|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2},$ (80)
Proof.
The proof strategy mirrors that of [23, Lemma V.18]. We utilize the perturbed insertion process $\check{Z}$ and define $U(X^{n},A^{n},Z^{n})\in\{t,0,1\}^{|{Y}|}$ for each bit in the output of size $|Y|$. Sequentially, for each bit in $Y$, we define the elements of $U$ based on the corresponding bit in $\check{Y}$ and the perturbed insertion process as follows:
  • For each bit present in both $Y$ and $\check{Y}$, the corresponding element of $U$ is defined as $t$. This case represents the bits originally existing in input $X$ and the insertions with the perturbed insertion process.
  • For each bit present in $Y$ but not in $\check{Y}$, we assign the corresponding bit in $U$ as either ‘0’ or ‘1’ based on the corresponding bit in $Y$. This case represents the reversed insertions due to the difference between the original insertion process and the perturbed insertion process.
Clearly, we have
$(X^{n},Y)$ $\mathrel{\mathop{\longleftrightarrow}\limits^{U}}(X^{n},\check{Y})$
$(X^{n},Y,K)$–→$\mathrel{\mathop{\leftarrow\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}\limits^{(U,Z)}}(X^{n},\check{Y},\check{K}).$ (81)
It is evident that we have:
$|H(\check{K}(X^{n})|X^{n},\check{Y}(X^{n}))-H$ $({K}(X^{n})|X^{n},{Y}(X^{n}))|\leq 2H(U)+H(Z).$ (82)
With the perturbed process, we further define $\check{z}=P(\check{Z}_{j}=1)$ for each $j$. The number of insertions reversed in a random run is at most
$\alpha^{2}\sum_{l_{0},l_{1}}p_{{L}}(l_{0},l_{1})(l_{0}+l_{1})^{2},$ (83)
in expectation, which can be bounded above by $4\alpha^{2}\mathbb{E}[{L}^{2}]$.
As in [23, Lemma V.3], there exists $\alpha_{0}$ such that the following occurs: Consider any $\gamma>1/2$, and define $\ell\equiv\lfloor 2\gamma\log(1/\alpha)\rfloor$. For all $\alpha<\alpha_{0}$, if $\mathbb{X}\in\mathcal{S}$ is such that $H(\mathbb{X})>1-\alpha^{\gamma}$, we have
$\sum_{l=\ell}^{\infty}lp_{L}(l)\leq 20\alpha^{\gamma}.$ (84)
Using [23, Lemma V.3] and ${L}\leq 1/\alpha$ w.p. 1, we have
$\mathbb{E}[{L}^{2}]\leq\frac{1}{\alpha}E[L]\leq 20\alpha^{\gamma-1},$ (85)
which implies
$\check{z}\leq 80\alpha^{1+\gamma}.$ (86)
It follows that
[…]
for small enough $\alpha$.
We further have $u\equiv P(U_{j}\neq t)$ for all $j$, where $u=\frac{\check{z}}{1+\alpha}\leq\check{z}$. It follows that $H(\mathbb{U})\leq u+h(u)\leq\alpha^{1+\gamma-\epsilon/2}$ for small enough $\alpha$. Using $H(\check{\mathbb{Z}})$ and $H(\mathbb{U})$, we have
$\lim_{n\rightarrow\infty}\frac{2H(U)+H(\check{Z})}{n}$ $=2(1+\alpha)H(\mathbb{U})+H(\check{\mathbb{Z}})$
$\leq 4H(\mathbb{U})+H(\check{\mathbb{Z}}),$ (88)
where the inequality is due to $\alpha\in[0,1]$. Hence, we get
$\lim_{n\rightarrow\infty}\frac{2H(U)+H(\check{Z})}{n}\leq 4\alpha^{1+\gamma-\epsilon/2},$ (89)
which completes the proof.∎
IV-D Calculations for $H(Y|X)$
Using the results given in Sections IV-A, IV-B and IV-C, we obtain the following corollary.
Corollary 1.
For an input process that is stationary and ergodic satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have
$\lim_{n\rightarrow\infty}$ $\frac{1}{n}H(Y|X^{n})$
$=h(\alpha)+\alpha\bigg{(}1-\frac{1}{2}\mathbb{E}[\log L_{0}]-\frac{1}{2}\sum_{l_{j}=1}^{\infty}\sum_{l_{j+1}=1}^{\infty}(l_{j+1}+1)$
$\qquad\cdot 2^{-l_{j}}2^{-l_{j+1}}h\big{(}\frac{1}{l_{j+1}+1}\big{)}\bigg{)}+\zeta,$ (90)
where $\zeta=\eta+\epsilon_{1}$ with
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v)\leq\eta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v),$
with $\epsilon>0$.
Proof.
We have
$H(Y|X^{n})$ $=H(Y,K|X^{n})-H(K|X^{n},Y)$ (91)
$=H(A^{n},B^{n})-H(A^{n},B^{n}|X^{n},Y,K)-H(K|X^{n},Y).$ (92)
Using Lemmas 4, 5 and 7 completes the proof. ∎
IV-E Achiveability
The next lemma proves the achievability of the described insertion channel model, which is obtained by using a Bernoulli i.i.d. input process.
Lemma 10 (Achievability).
Let $\mathbb{X}^{*}$ be the iid Bernoulli(1/2) process. For any $\epsilon>0$, we have
$C(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}).$ (93)
Proof.
Since $\mathbb{X}$ is i.i.d. Bernoulli(1/2) with size $n$, it has a run-length distribution $p_{L}(l)=2^{-l}$. Additionally, the corresponding output $Y(X^{*,n})$ is also i.i.d. Bernoulli with length $L=n+\text{Bin}(n,\alpha)$. Hence, we have
$H(Y,L)$ $=H(L)+H(Y|L)$
$=H(L)+\sum_{l=n}^{2n}H(Y|L=l)p_{L}(l)$
$=H(L)+\sum_{l=n}^{2n}l2^{-l}$
$=\mathcal{O}(\log(n))+n(1+\alpha).$ (94)
As a result,
$H(Y(X^{*,n}))$ $=\mathcal{O}(\log(n))+n(1+\alpha).$ (95)
Using the estimate of $H(Y|X^{*,n})$ from the Corollary 1 with $P(z=a,v=b)=\mathcal{O}(\alpha^{2})$ and $\mathbb{E}[L_{0}\log L_{0}]<\infty$ completes the proof. ∎
IV-F Converse
We denote the processes with maximum run length $L$ with probability one by $\mathcal{S}_{L}$. Without loss of generality, we restrict ourselves to finite-length runs. With the next lemma, we illustrate that restricting ourselves to $\mathcal{S}_{L^{*}}$ for large enough $L^{*}$ will not cause a significant loss.
Lemma 11.
For any $\epsilon>0$ there exists $\alpha_{0}=\alpha_{0}(\epsilon)>0$ such that the following happens for all $\alpha<\alpha_{0}$. For any $\mathbb{X}\in\mathcal{S}$ such that $H(\mathbb{X})>1+2\alpha\log\alpha$ and for any $L^{*}>\log(1/\alpha)$, there exists $\mathbb{X}_{L^{*}}\in\mathcal{S}_{L^{*}}$ such that
$I(\mathbb{X})\leq I(\mathbb{X}_{L^{*}})+\alpha^{1/2-\epsilon}(L^{*})^{-1}\log(L^{*}).$ (96)
Proof.
We construct $\mathbb{X}_{L^{*}}$ by flipping a bit each time it is the $(\mathcal{S}_{L^{*}}+1)$-th consecutive bit with the same value. The density of such bits is upper bounded by $\beta=\frac{P(L_{0}>{L^{*}})}{L^{*}}$. The corresponding output is denoted by $Y_{L^{*}}=Y(X^{n}_{L^{*}})$.
We further define $F=F(\mathbb{X},\mathbb{D})$ as a binary vector with the same length as the channel output. The elements of $F$ are 1 wherever the corresponding bit in $Y_{L^{*}}$ is flipped relative to $Y$, and 0 otherwise. The entropy $H(F)$ can be calculated by:
$H(F)$ $=H(F|L_{F})+H(L_{F})$ (97)
$\leq H(F|L_{F})+\log(n+1)$ (98)
$=\sum_{l=n}^{2n}H(F|L_{F}=l)P(L_{F}=l)+\log(n+1)$ (99)
$=\sum_{l=n}^{2n}h(\beta)lP(L_{F}=l)+\log(n+1)$ (100)
$=h(\beta)(1+\alpha)n+\log(n+1),$ (101)
where $L_{F}$ is the length of $F$. The rest of the proof follows the same steps as [22, Lemma III.2]. ∎
Lemma 12 (Converse).
For any $\epsilon>0$ there exists there exists $\alpha_{0}=\alpha_{0}(\epsilon)>0$ such that the following happens. For any $L^{*}\in\mathbb{N}$ and any $\mathbb{X}\in\mathcal{S}_{L^{*}}$ if $\alpha<\alpha_{0}(\epsilon)$, then
$I(\mathbb{X})\leq 1+\alpha\log\alpha+G_{1}\alpha+\alpha^{2-\epsilon}(1+\alpha^{1/2}L^{*}).$ (102)
Proof.
We can write
$H(Y)\leq n(1+\alpha)+\log(n+1),$ (103)
since $Y(X^{n})$ contains $n+\text{Binomial}(n,\alpha)$ bits. The proof involves using the lower bound in Corollary 1 along with $h(z,v)$. Note that for $H(\mathbf{X})>1+2\alpha\log\alpha$, it is also true that $H(\mathbf{X})>1-\alpha^{1-\epsilon}$. Hence Corollary 1 with $\alpha=1-\epsilon$ applies. Furthermore, recall that the joint PMF of $Z$ and $V$ was given in (36). One can upper bound the probabilities as
$P(z=0,v=0)$ $\leq 1-2\alpha^{2}\mathbb{E}[L_{0}],$ (104)
$P(z=1,v=0)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}],$
$P(z=1,v=1)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}].$
We employ [22, Lemma IV.3], which gives $\mathbb{E}[L_{0}]\leq K_{1}(1+\sqrt{\alpha(\log(1/\alpha))^{3}L^{*}})$; hence, we obtain
$h(z,v)\leq 0.5\alpha^{2-\epsilon}(2+0.5\alpha^{1/2}L^{*}),$ (105)
for all $\alpha\leq\alpha_{0}$ with $\alpha_{0}=\alpha_{0}(\epsilon)>0$. Also noting
$\big{|}\mathbb{E}[\log L_{0}]-\sum_{l=1}^{\infty}2^{-l-1}l\log l\big{|}=o(\alpha^{1/2-\epsilon}\log L^{*}).$ (106)
The proof follows. ∎
V Capacity Approximation for the Gallager Insertion Channel
In this section, we provide the capacity approximation for the Gallager insertion channel model, similar to the technique used for the first insertion channel model. The main result for the capacity is presented in the following theorem.
Theorem 13.
Let $C(\alpha)$ be the capacity of the insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C(\alpha)=1+\alpha\log(\alpha)+G_{2}\alpha+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{O}(\alpha^{3/2-\epsilon})}},$ (107)
where
$G_{2}=-$ $\log(e)-\frac{7}{8}+\frac{1}{4}\sum_{l=1}^{\infty}2^{-l-1}l\log l$
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\frac{1}{4}}}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right),$ (108)
which is approximated as $G_{2}\approx-0.5865$.
In the next subsections, we first provide the different terms needed to compute $I(X^{n},Y(X^{n}))$, then show the achievability and converse proofs leading to this approximation.
V-A Calculation of $H(A^{n},B^{n})$
We follow a similar approach to that used for the simple insertion channel model as described in Sec. IV-A. The insertion pattern process corresponding to the sequence of insertion channels is defined as $\mathbb{A}$, an i.i.d. Bernoulli($\alpha$) process that determines the locations of insertions, independent of the input, and represented by the $n$-bit binary vector $A^{n}$. Each element of $A^{n}$, denoted as $A_{i}$, is set to 1’ if the $i$-th bit of the input vector is replaced by two random bits, and 0’ otherwise. For the inserted bits, as described in Section IV-A, we use the notation $B_{i}^{j}$ for $j\in\{1,2\}$.
Lemma 14.
We have
$\lim_{n\rightarrow\infty}\frac{1}{n}H(A^{n},B^{n})=h(\alpha)+2\alpha,$ (109)
where $h(\cdot)$ is the binary entropy function.
Proof.
Note that for this channel model, we have $B^{n}=(B^{1,n},B^{2,n})$.
The joint probability $P(A_{i}=a,B^{1,n}_{i}=b^{1},B^{2,n}_{i}=b^{2})$ can be expressed as:
$P(a,b^{1},b^{2})=\begin{cases}\frac{\alpha}{4},&\text{if }(a,b^{1},b^{2})\in\{(1,1,0),(1,0,1),\\ &\qquad\qquad(1,1,1),(1,0,0)\}\\ 1-\alpha,&\text{if }(a,b^{1},b^{2})=(0,0,0)\\ 0,&\text{otherwise}\end{cases}$
Utilizing this joint probability, we can derive the following entropy:
$\frac{1}{n}H(A^{n},B^{1},B^{2})=-\alpha\log$ $\alpha+2\alpha-(1-\alpha)\log(1-\alpha).$ (110)
V-B Calculating $H(A^{n},B^{n}|X^{n},Y,K)$
As in Section IV-B, we introduce a modified insertion process for the Gallager insertion channel by defining $\hat{A}^{n}$ and $\hat{B}^{n}=(\hat{B}^{1,n},\hat{B}^{2,n})$. For the extended runs, the modified process keeps the runs with at most one insertion as they are while reversing more insertions.
We further define
$\mathbb{Z}^{n}=\hat{\mathbb{A}}^{n}\oplus\mathbb{A}^{n},$ (111)
and
$\mathbb{V}^{n}=\hat{\mathbb{B}}^{n}\oplus\mathbb{B}^{n},$ (112)
where $\mathbb{V}^{n}=(\mathbb{V}^{1,n},\mathbb{V}^{2,n})$ and can be rewritten as
$\mathbb{V}^{1,n}$ $=\hat{\mathbb{B}}^{1,n}\oplus\mathbb{B}^{1,n},$
$\mathbb{V}^{2,n}$ $=\hat{\mathbb{B}}^{2,n}\oplus\mathbb{B}^{2,n}.$ (113)
As in the first insertion channel model, we use the following pairs to show that the modified insertion process do not affect the first two order terms:
$\left((X^{n},Y,K,A^{n},B^{n}),(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n})\right),$ (114)
and
$\left((X^{n},Y,K),(X^{n},\hat{Y},\hat{K})\right).$ (115)
With a similar approach, both of these pairs are in the form of $(T,E)$ such that $T$ is a function of $(E,Z^{n},V^{n})$ and $E$ is a function of $(T,Z^{n},V^{n})$. Hence, we have:
$|H(T)-H(E)|\leq H(Z^{n},V^{n}),$ (116)
where ${V}^{n}=({V}^{1,n},{V}^{2,n})$. As a result, we have:
$|H($ $A^{n},B^{n}|X^{n},Y,K)-H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})|\leq 2H(Z^{n},{V}^{1,n},{V}^{2,n}).$ (117)
For these processes and to calculate $H(Z^{n},{V}^{1,n},{V}^{2,n})$, we have the following probabilities:
$P$ $(z=a,v^{1}=b,v^{2}=c)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}],\\ &\qquad\text{if }(a,b,c)=(0,0,0)\\ &\frac{1}{4}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}]\right),\\ &\qquad\text{if }(a,b,c)\in\mathcal{A}\\ &0,\qquad\text{otherwise},\end{cases}$ (118)
where $\mathcal{A}=\{(1,0,0),(1,0,1),(1,1,0),(1,1,1)\}$.
As in Sec. IV-B, one can use (118) to calculate $H(Z^{n},{V}^{1,n},{V}^{2,n})\leq nh(z,v^{1},v^{2})$ and show that it does not affect lower order terms. Comparing (118) with (36) (from the previous channel model), one can easily verify that $h(z,v^{1},v^{2})\leq h(z,v)$.
Lemma 15.
Consider the Gallager insertion channel model with any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}<\infty]$. Then, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}H({A}^{n},$ ${B}^{n}|X^{n},{Y}(X^{n}),K)$
$=\frac{1}{4}\alpha\mathbb{E}[\log(L_{0})]+\frac{1}{4}\alpha\frac{\mathbb{E}[L_{0}]-1}{\mathbb{E}[L_{0}]}-\eta,$ (119)
where $-2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2})$.
Proof.
The proof follows from (117) and Lemma 16 given in the next part. ∎
Lemma 16.
With Gallager insertion channel model, consider any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}<\infty]$. Then, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\hat{A}^{n},$ $\hat{B}^{n}|X^{n},\hat{Y}(X^{n}),\hat{K})$
$=\frac{1}{4}\alpha\mathbb{E}[\log(L_{0})]+\frac{1}{4}\alpha\frac{\mathbb{E}[L_{0}]-1}{\mathbb{E}[L_{0}]}-\delta,$ (120)
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)$.
Proof.
To simplify understanding of the lemma, let’s focus on a run of ‘0’s with length $l_{r}$, for example:
[…]
For such a run, there are five different cases to consider:
  1. 1.
    No insertion: In this case, there is no ambiguity and no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$ as discussed in Section IV-B.
  2. 2.
    Replace a bit with ‘11’: Inserting ‘11’ into the run of ‘0’s reveals all information regarding $\hat{D}$, $\hat{B}^{1,n}$, and $\hat{B}^{2,n}$ given the marker vector $\hat{K}$. Hence, there’is no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
  3. 3.
    Replace a bit with ‘00’: There are $l_{r}$ possible locations (each equally likely) for insertion in $\hat{D}$ given the marker vector $K$. No additional ambiguity arises from $\hat{B}^{n}$ (the corresponding bits of $\hat{B}^{i,n}$ are all zero regardless of the $l_{r}$ possible insertion positions for $i\in\{1,2\}$). The corresponding contribution is:
    $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})\Big{|}_{00}=\frac{1}{4}\sum_{i=1}^{n}\hat{A}_{i}\log(l_{(i)}),$ (122)
    where $l_{(i)}$ is the length of the run containing the $i$-th bit.
    Your explanation is clear, but I’ll make a few adjustments for presentation and clarity:
  4. 4.
    Replace a bit with ‘01’: For this case, we can consider two disjoint events as follows:
    • The insertions (replacements with ‘01’) for the first $l_{r}-1$ bits: Consider the run given in (121). Assume we replace the first bit with ‘01’:
[…]
      Now, consider another case in which the second bit is replaced by ‘10’:
[…]
      Clearly, these two cases result in the same output. Hence, there is a contribution of $\log 2$ for these first $l_{r}-1$ bits.
    • The insertion (replacement with ‘01’) for the last bit of the run: If the last bit is replaced with ‘01’, then there will be no ambiguity.
    We can write the contribution from this case as:
    $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})\Big{|}_{01}$ $=\frac{1}{4}\sum\limits_{\begin{subarray}{c}i=1,\\ i\text{ is not the}\\ \text{last bit of a run}\end{subarray}}^{n}\hat{A}_{i}\log(2).$ (125)
  5. 5.
    Replace a bit with ‘10’: Similar to the previous case, we have two disjoint events.
    • The insertions (replacements with ‘10’) for the last $l_{r}-1$ bits: Consider the run given in (121). Assume we replace the second bit with ‘10’:
[…]
      Now, consider another case in which the first bit is replaced by ‘01’:
[…]
      Clearly, these two cases result in the same output. Hence, there is a contribution of $\log 2$ for these last $l_{r}-1$ bits. We further note that this event intersects the scenario we described previously with the replacement with ‘01’. Hence, we will discount their intersection while calculating the overall ambiguity.
    • The insertion (replacement with ‘10’) for the first bit of the run: If the first bit is replaced with ‘10’, then there will be no ambiguity.
    We can write the contribution from this case as:
    $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})\Big{|}_{01}$ $=\frac{1}{4}\sum\limits_{\begin{subarray}{c}i=1,\\ i\text{ is not the}\\ \text{first bit of a run}\end{subarray}}^{n}\hat{A}_{i}\log(2).$ (128)
Hence, considering the insertions with ambiguity (and the intersecting events of insertions with ‘01’ and ‘10’), the overall contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$ can be written as:
$H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})=\frac{1}{4}$ $\sum_{i=1}^{n}\hat{A}_{i}\log(l_{(i)})+\frac{1}{4}\sum\limits_{\begin{subarray}{c}i=1,\\ i\text{ is not the}\\ \text{first bit of a run}\end{subarray}}^{n}\hat{A}_{i}\log(2).$ (129)
To account for the first bit of the run, (129) can be rewritten as:
$H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})=\frac{1}{4}$ $\sum_{i=1}^{n}\hat{A}_{i}\log(l_{(i)})+\frac{1}{4}\sum\limits_{i=1}^{n}\hat{A}_{i}\log(2)\frac{l_{(i)}-1}{l_{(i)}}.$ (130)
We will use the same definition as in (48):
$P(\hat{A}_{i}=1)\in(\alpha-2l_{(i)}\alpha^{2},\alpha).$ (131)
Inserting (131) into (130), taking expectation and letting $n\rightarrow\infty$, we have
$H($ $\hat{A}^{n},\hat{B}^{1},\hat{B}^{2}|X^{n},\hat{Y},\hat{K})=\frac{1}{4}n\alpha\mathbb{E}[\log(L_{0})]+\frac{1}{4}n\alpha\frac{\mathbb{E}[L_{0}]-1}{\mathbb{E}[L_{0}]}-\delta n,$ (132)
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)$, which completes the proof of Lemma 16. ∎
V-C Calculating $H(K|X^{n},Y)$
Lemma 17.
For the Gallager insertion channel with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(K|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)+\epsilon_{1},$ (133)
where
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (134)
with $\epsilon>0$.
Proof.
To compute $H(K|X^{n},Y)$, we define the perturbed insertion process $\check{\mathbb{Z}}$ as in Section IV-C, and follow the same steps. We use the same definitions $\check{K}$, $\check{Y}$, and $\bar{Y}$. With this process, we assume that there is at most one insertion on consecutive two runs.
We compute $H(\check{K}|X^{n},\check{Y})$ and find a bound for $|H({K}|X^{n},{Y})-H(\check{K}|X^{n},\check{Y})|$, as in the following Lemmas 18 and 19, respectively. ∎
Lemma 18.
For the Gallager insertion channel with the perturbed insertion process and a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(\check{K}|X^{n},\check{Y})=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}+\epsilon_{3}$ (135)
which is approximately $\sim 1.4090\alpha-\epsilon_{2}+\epsilon_{3}$ where
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}$ $\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right),$ (136)
and
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (137)
Proof.
To calculate $H(\check{K}|X^{n},\check{Y})$ as in (IV-C), the first step is to calculate
$t_{j}\equiv H\left(|\check{Y}(j)|\Big{|}X(j)\cdots X(M),\bar{Y}(j^{\prime})\cdots\bar{Y}(M^{\prime})\right).$ (138)
We define the length of $\bar{Y}(j)$ by $L_{j}$, i.e., $|\bar{Y}(j)|=L_{j}$. We have four cases:
  1. 1.
    Case $\mathcal{V}_{1}$: $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: In this case, we certainly do not have insertion in $X(j)$. Hence, $t_{j}=0$.
  2. 2.
    Case $\mathcal{V}_{2}$: $|\bar{Y}(j)|<|X(j)|$: We have a single bit replacement inside $X(j)$. Consider $X(j)$ as a run of 0’s:
    • There can be a replacement with ‘11’ or ‘10’ (on any bit location).
    • There can be a replacement with ‘01’ (except the last bit).
    In those cases, there is no ambiguity in the dimension of $Y(j)$ (or equivalently on $\check{K}$). Hence, $t_{j}=0$.
  3. 3.
    Case $\mathcal{V}_{3}$: $|\bar{Y}(j)|=|X(j)|+1$, $|\bar{Y}(j+1)|=|X(j+1)|$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs (since there will be at most a single insertion in two consecutive runs, we know that there is no insertion at the $(j+2)$-th run).
    Consider the following example:
    $X(j)X(j+1)$ =[0 0 0 1 1 1 1],absentdelimited-[]$0$\ 0\ 1\ 1\ 1\ 1],= [ 0 0 0 1 1 1 1 ] , (139)
    $\bar{Y}(j)\bar{Y}(j+1)$ =[0 0 0 0 1 1 1 1].absentdelimited-[]$0$\ 0\ 0\ 1\ 1\ 1\ 1].= [ 0 0 0 0 1 1 1 1 ] . (140)
    This can happen in two different scenarios:
    - If any of the bits in $X(j)$ are replaced by ‘00’, which is with probability:
    $P_{a}=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}|X(j)|.$ (141)
    - Or, if the first bit of $X(j+1)$ is replaced by ‘01’, which is with probability:
    $P_{b}=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}.$ (142)
    Note that these two scenarios are disjoint. Hence, we have
    $P(\mathcal{V}_{3})=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}(|X(j)|+1).$ (143)
  4. 4.
    Case $\mathcal{V}_{4}$: $|\bar{Y}(j)|=|X(j)|$, $|\bar{Y}(j+1)|=|X(j+1)|+1$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs.
    Consider the following example:
    $X(j)X(j+1)$ =[0 0 0 1 1 1 1],absentdelimited-[]$0$\ 0\ 1\ 1\ 1\ 1],= [ 0 0 0 1 1 1 1 ] , (144)
    $\bar{Y}(j)\bar{Y}(j+1)$ =[0 0 0 1 1 1 1 1].absentdelimited-[]$0$\ 0\ 1\ 1\ 1\ 1\ 1].= [ 0 0 0 1 1 1 1 1 ] . (145)
    This can happen in two different scenarios:
    - If any of the bits in $X(j+1)$ are replaced by ‘11’, which is with probability:
    $P_{e}=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}|X(j+1)|.$ (146)
    - Or, if the last bit of $X(j)$ is replaced by ‘01’, which is with probability:
    $P_{f}=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}.$ (147)
    Using these disjoint events, we have
    $P(\mathcal{V}_{4})=\frac{1}{4}\alpha(1-\alpha)^{|X(j)|+|X(j+1)|-1}(|X(j+1)|+1).$ (148)
Overall, the only ambiguity in $\check{K}$ comes from two cases: for $|\check{Y}(j)|=|X(j)|$ and $|\check{Y}(j)|=|X(j)|+1$. Hence, by defining $r_{j}\equiv|X(j)|$, we can write
$|\check{Y}(j)|=\begin{cases}r_{j}+1,&\text{wp }\frac{1}{4}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}(r_{j}+1)\\ r_{j},&\text{wp }\frac{1}{4}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}(r_{j+1}+1)\end{cases}$ (149)
Since the only ambiguity on $|\check{Y}(j)|$ comes from (149), we can normalize it as:
$|\check{Y}(j)|=\begin{cases}r_{j}+1,&\text{wp }\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\\ r_{j},&\text{wp }\frac{r_{j+1}+1}{r_{j}+r_{j+1}+2}\end{cases}$ (150)
As a result, we have
$t_{j}=h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right).$ (151)
Thus, the expected contribution of this term to the sum is
$\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}\frac{1}{4}\alpha(1-\alpha)^{r_{j}+r_{j+1}-1}(r_{j}+r_{j+1}+2)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right),$ (152)
where $p_{L(k)}(l_{1},\dots,l_{k})$ is the joint distribution of $k$ input run lengths.
Note that we have $(1-\alpha)^{r_{j}+r_{j+1}-1}\in\left(1-\alpha(r_{j}+r_{j+1}),1\right)$, which gives
$H(\check{K}|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}$ (153)
with
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right).$ (154)
where $p_{L(2)}(r_{j},r_{j+1})\leq 1$ is used in (154).
As in Lemma 8, to bound (153), let us consider $p^{*}_{L(k)}(l_{1},\dots,l_{k})=2^{-\sum_{i=1}^{k}l_{i}}$ and $H(\mathbb{X})>1-\alpha^{\gamma}$ with $\gamma>1/2$. Then, by applying (76) to the first term in (153),
$H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (155)
with
$A_{1}=\frac{1}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)$ (156)
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (157)
and $\epsilon_{2}$’s range is defined in (154).
Note that we have $A_{1}\approx 1.4090$. ∎
Lemma 19.
With a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$; the difference between $H(K|X^{n},Y)$ and $H(\check{K}|X^{n},\check{Y})$, corresponding to the original insertion and perturbed insertion processes with the Gallager insertion channel, respectively, satisfies:
$|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2}.$ (158)
Proof.
The proof is similar to that of Lemma 9. ∎
V-D Calculations for $H(Y|X)$
Using the previously mentioned results in Sections IV-A, IV-B and IV-C, we obtain the following corollary.
Corollary 2.
For the Gallager insertion channel model with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$\lim_{n\rightarrow\infty}$ $\frac{1}{n}H(Y|X^{n})$
$=h(\alpha)+\alpha\Bigg{(}2-\frac{1}{4}\mathbb{E}[\log(L_{0})]-\frac{1}{4}\frac{\mathbb{E}[L_{0}]-1}{\mathbb{E}[L_{0}]}$
$\qquad-\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)\Bigg{)}$
$\qquad+\zeta,$ (159)
where $\zeta=\eta+\epsilon_{1}$ with
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2}),$
with $\epsilon>0$.
Proof.
We have
$H(Y|X^{n})$ $=H(A^{n},B^{n})-H(A^{n},B^{n}|X^{n},Y,K)-H(K|X^{n},Y).$ (160)
Using Lemmas 14, 15, and 17 completes the proof. ∎
V-E Achievability and Converse
The achievability and converse proofs follow a similar structure to those of the first insertion channel model (see Sections IV-E and IV-F). Note that for the converse proof, when calculating $h(z,v^{1},v^{2})$, we observe from (36) and (118) that the bound from (105) can be applied, as $h(z,v^{1},v^{2})\leq h(z,v)$.
VI Conclusions
In this paper, we focused on the capacity of insertion channels with small insertion probabilities, a topic that has been largely overlooked in the literature despite its relevance in practical applications such as DNA coding and storage. Our approach builds upon the series expansion of a previously known capacity point, specifically a channel with no insertions (i.e., zero insertion probability), where the capacity is 1 and is achieved with an i.i.d. Bernoulli$(1/2)$ input distribution. By leveraging this known point, we expect that for small insertion probabilities, the achievable rate will not deviate significantly from this value. With this intuition, we derived the first two terms of the capacity expansion by decomposing the expressions necessary to achieve the maximum rate. Using this methodology, we obtained capacity approximations for two fundamental insertion channel models: 1) the insertion channel with random bit insertions, and 2) the Gallager insertion channel, where a bit is replaced by two random bits. Our results demonstrate that these approximations are achievable by a uniform i.i.d. input distribution up to the first two orders, as expected. For future work, possible extensions include applying similar methods to different channel models or symbol alphabets. For instance, the binary input case could be extended to nonbinary inputs, e.g., 4-ary for DNA storage systems, with deletion and/or insertion channels. Further research could also consider different channel models, such as the absorption channel or the insertion, deletion, and substitution (IDS) channel. Thus, our approach lays the foundation for various extensions and can facilitate capacity analysis for these different models.